A 5cm² cord has original length of 120cm is pulled by a 20N force. The change in length in the cord is 5mm. Find the strain, stress and elastic modulus.
1
Expert's answer
2020-10-30T13:19:30-0400

By definition, the stress is the ratio of a force to a cross-sectional area:


"\sigma = \dfrac{F}{A} = \dfrac{20N}{5\times 10^{-4}m^2} = 4\times 10^4Pa"

The strain is the ratio of a length change to the original length:


"\varepsilon = \dfrac{\Delta L}{L} = \dfrac{0.5cm}{120cm} \approx 4.17\times 10^{-3}"

The Hook's law states that:


"\sigma = E\varepsilon"

where "E" is the elastic modulus. Thus:


"E = \dfrac{\sigma}{\varepsilon} = \dfrac{4\times 10^4Pa}{4.17\times 10^{-3}} = 9.6\times 10^6Pa = 9.6\space GPa"

Answer. stress = "4\times 10^4Pa", strain = "4.17\times 10^{-3}", elastic modulus = "9.6\times 10^6Pa".